OpenStudy (anonymous):

compute the monthly add on interest loan of $1290 with an annual interest rate of 7 percent and a term of two years

OpenStudy (anonymous):

I you are talking about compound interest obviously as the interest is aded on every month to give a new principal then the interest every month will change.For example Principal yr 1: $1290 Interest yr 1(7%p/m.): $90.60 Principal yr 2: $1380.30 [because interest is added on] Interest yr 2(7% p/m.): $96.60 However, it may be possible to find compound interest over 24 months (2 years) an then get an average per month as follows: Pn = $1290(1 + 7/100)^23 Pn = $1290(1.07)^23 Pn = $1290(4.74) Pn = $6115.28 Compound interest is therefore: $6115.28 - $1290 = $4825.28 divide this by 24 months (2 years) to get an average monthly interest: $4825.28/24 = $201.25